A circuit is of a 50 W heater connected to a four-cell battery. a) Calculate the potential difference of the battery. (Each cell is 3.5 V). b) Calculate the flowing through the heater. c) Write down the equation that links energy transferred, power and time. d) Calculate the energy transferred to the heater when it is used for 10 minutes.
1
Expert's answer
2018-04-24T08:51:08-0400
The potential difference of the battery
E= E_n·n=3.5·4=14 V
I=P/E=50/14=3.57 W
W=P·t , where t measured in hours
The energy transferred to the heater when it is used for 10 minutes
W=50·1/6=8.33 W·h